Consider the following information on factor endowments for the two countries of Spain and Germany. Answer the questions following the table on the discussion board. Answer each question separately by the number.

Factor Endowments Spain Germany
Labor Force 30 million 15 million
Capital Stock 200,000 machines 400,000 machines
  1. Which country is relatively labor abundant? Show how you know.
  2. Which country is relatively capital abundant? Show how you determine.
  3. Assuming that steel is relatively capital intensive production, which country has a comparative advantage in production of steel? Assuming apparel is relatively labor intensive production, which country has the comparative advantage in production of apparel? Explain your answers.
  4. If these two countries agrees to trade, what kind of restructuring would the countries face? Would there be complete specialization in each country? Would capital and labor remain fully employed in both the countries with persisting trade? Why or why not?
  5. How would income distribution for labor and capital change in each of the two countries with persisting trade? Explain your answer.
Answer #1

Answers:

1. Which country is relatively labor abundant? Spain..

For Spain: Labor Force/Capital= 30,000,000/2,00,000 = 150.

For Germany: Labor Force/Capital= 15,000,000/4,00,000 = 37.5.

2. Which country is relatively capital abundant? Germany.

For Spain: Capital Stock/Labor Force= 2,00,000/30,000,000 = 0.006

For Germany: Capital Stock/Labor Force= 4,00,000/15,000,000 = 0.026

3. Assuming that steel is relatively capital intensive production, which country has a comparative advantage in production of steel? Germany.

4. Assuming apparel is relatively labor intensive production, which country has the comparative advantage in production of apparel? Explain your answers. Spain.

Explanation for Ans 3&4: According to Heckscher–Ohlin model (H–O model) theory, A country will produce more good that makes the use of the factor in which it is abundant. That is if any country has more capital it will deal in capital intensive production and if any country has more labor it will deal in labor intensive production.